OpenStudy (anonymous):

this is my last question:)) tilll im done with my assignment! please help! Write a system of two equations with the variables m and f in each to represent the following problem: To skate at Roller Heaven, each person must pay a monthly membership fee (m) and a fee for each session (f). In January, Art attended 5 sessions and paid $25 in all. In July, he attended 10 sessions, and he paid $40 in all.

OpenStudy (hoblos):

m + 5f = 25 m + 10f = 40 -5f = -15 => f=3 m=25-15 =10